OpenStudy (anonymous):

(xlny+xy) if i differentiate that with respect to y... i think i get 0

OpenStudy (anonymous):

When you differentiate with respect to y, all you do is treat the \(x\) variable as a constant. Since \(\ln{y}\) doesn't differentiate to zero, and neither does \(y\), you shouldn't get zero.

OpenStudy (anonymous):

ok wat do i get

OpenStudy (anonymous):

cn you show me

OpenStudy (anonymous):

how to do it den

OpenStudy (anonymous):

So the derivative of \(\ln{y}\) is \(\frac{1}{y}\), and the derivative of \(y\) is \(1\). Does that help get you moving?

myininaya (myininaya):

He didn't say partial derivative

OpenStudy (anonymous):

i got the answer x/y+x

myininaya (myininaya):

This could be implicit differiention

OpenStudy (anonymous):

is that rite

OpenStudy (anonymous):

Yep, that's right. Unless myininaya is right about it being implicit differentiation. Are you learning about implicit differentiation now, or partial derivatives?

myininaya (myininaya):

If we aren't doing partial differentiation, the answer is x'lny+x/y+x'y+x

OpenStudy (anonymous):

implicit

myininaya (myininaya):

Since x is a function of y

myininaya (myininaya):

Have you heard of partial derivatives?